OpenStudy (samanddean1967):

Area of an isosceles triangle is 12 ft squared, the base is 6, what is the leg of the triangle Im not sure what it is asking @kittiwitti

OpenStudy (samanddean1967):

@kittiwitti1

OpenStudy (samanddean1967):

the leg of the triangle would be the base right?

Directrix (directrix):

A = 1/2 * 6 * h 12 = 1/2 * 6 * h h = ? Get h first and then find the leg.